# Overcoming label shift in targeted federated learning

### v1

@@ 3.2 FedPALS in the limits
Now, consider when the target label distribution is in the convex hull of the source label distributions, $\mathrm{Conv}(S)$. Then, we can find a convex combination $\alpha^{c}$ of source distributions $S_{i}(Y)$ that recreate $T(Y)$, that is, $T(Y)=\sum_{i=1}^{M}\alpha_{i}^{c}S_{i}(Y)$. However, when there are more clients than labels, $M>K$, such a satisfying combination $\alpha^{c}$ need not be unique and different combinations may have different effective sample size. Let $A^{c}=\{\alpha^{c}\in\Delta^{M-1}:T(Y)=\alpha^{c}S(Y)\}$ denote all satisfying combinations. For a sufficiently small regularization penalty $\lambda$, the solution to (7) will be the satisfying combination with largest effective sample size.
@@
If there are fewer clients than labels, $M<K$, the set of target distributions for which a satisfying combination exists has measure zero, see Figure 1 (left). Nevertheless, the two cases above allow us to interpolate between being as faithful as possible to the target label distribution ($\lambda\to 0$) and retaining the largest effective sample size ($\lambda\to\infty$), the latter coinciding with FedAvg.
Finally, when $T\in\mathrm{Conv}(S)$ and $\lambda\rightarrow 0$, Proposition 1 applies also to FedPALS; the aggregation strategy results in an unbiased estimate of the target risk gradient in the stochastic gradient descent (SGD) setting. This does not necessarily hold when clients perform multiple local updates before aggregation. In practice, many federated learning systems, including FedAvg, allow clients to perform several steps of local optimization (e.g., over an entire epoch) before aggregating the parameter updates at the server. This is a limitation of our theoretical understanding: when clients are not constrained to a single SGD step, the aggregated updates no longer strictly minimize the target risk.
@@
If the target distribution does not lie in $\mathrm{Conv}(S)$, see Figure 1 (left), FedPALS projects the target to the “closest point” in $\mathrm{Conv}(S)$ if $\lambda=0$, and to a tradeoff between this projection and the FedAvg aggregation if $\lambda>0$.
Choice of hyperparameter $\lambda$
A salient question in Cases 2 & 3 is how to choose the strength of the regularization, $\lambda$. A larger value will generally favor influence from more clients, provided that they have sufficiently many samples. When $T\not\in\mathrm{Conv}(S)$, the convex combination closest to $T$ could have weight on a single vertex. This will likely hurt the generalizability of the resulting classifier. In experiments, we compare values of $\lambda$ that yield different effective sample sizes, such as 10%, 25%, 50%, 75%, or 100% of the original sample size, $N$. We can find these using binary search by solving (7) and calculate the ESS. One could select $\lambda$ heuristically based on the the ESS, or treat $\lambda$ as a hyperparameter and select it using a validation set. Although this would entail training and evaluating several models which can be seen as a limitation.
@@
In problems with a large number of labels, $K\gg 1$, it is common that any individual domain (clients or target) supports only a subset of the labels. For example, in the IWildCam benchmark, not every wildlife camera captures images of all animal species. When the target $T(Y)$ is sparse, meaning $T(y)=0$ for certain labels $y$, it becomes easier to find a good match $\alpha^{\lambda}S(Y)\approx T(Y)$ if the client label distributions are also sparse. Achieving a perfect match, i.e., $T\in\mathrm{Conv}(S)$, requires that (i) the clients collectively cover all labels in the target, and (ii) each client contains only labels that are present in the target. If this is also beneficial for learning, it would suggest that the client-presence of labels that are not present in the target would harm the aggregated model. We study the implications of sparsity of label distributions empirically in Section 5.
@@ 4 Related work
Efforts to mitigate the effects of distributional shifts in federated learning can be broadly categorized into client-side and server-side approaches. Client-side methods use techniques such as clustering clients with similar data distributions and training separate models for each cluster (Ghosh et al., 2020; Sattler et al., 2020; Vardhan et al., 2024), and meta-learning to enable models to quickly adapt to new data distributions with minimal updates (Chen et al., 2018; Jiang et al., 2019; Fallah et al., 2020b). Other notable strategies include logit calibration (Zhang et al., 2022), regularization techniques that penalize large deviations in client updates to ensure stable convergence (Li et al., 2020b, 2021), and recent work on optimizing for flatter minima to enhance model robustness (Qu et al., 2022; Caldarola et al., 2022). Server-side methods focus on improving model aggregation or applying post-aggregation adjustments. These include optimizing aggregation weights (Reddi et al., 2021), learning adaptive weights (Li et al., 2023), iterative moving averages to refine the global model (Zhou et al., 2023), and promoting gradient diversity during updates (Zeng et al., 2023). Both categories of work overlook shifts in the target distribution, leaving this area unexplored.
Another related area is personalized federated learning, which focuses on fine-tuning models to optimize performance on each client’s specific local data (Collins et al., 2022; Boroujeni et al., 2024; Fallah et al., 2020a). This setting differs fundamentally from our work, which focuses on improving generalization to new target clients without any training data available for fine-tuning. Label distribution shifts have also been explored with methods such as logit calibration (Zhang et al., 2022; Wang et al., 2023; Xu et al., 2023), novel loss functions (Wang et al., 2021), feature augmentation (Xia et al., 2023), gradient reweighting (Xiao et al., 2023), and contrastive learning (Wu et al., 2023). However, like methods aimed at mitigating the effects of general shifts, these do not address the challenge of aligning models with an unseen target distribution, as required in our setting.
Generalization under domain shift in federated learning remains underdeveloped (Bai et al., 2024). One notable approach is tailored for medical image segmentation, where clients share data in the frequency domain to achieve better generalization across domains (Liu et al., 2021). However, this technique requires data sharing, making it unsuitable for privacy-sensitive applications like ours. A different line of work focuses on addressing covariate shift in federated learning through importance weighting (Ramezani-Kebrya et al., 2023). Although effective, this method requires sending samples from the test distribution to the server, which violates our privacy constraints.
@@ 5 Experiments
While numerous benchmarks exist for federated learning (Caldas et al., 2018; Ogier du Terrail et al., 2022; Chen et al., 2022) and domain generalization (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021), respectively, until recently none have addressed tasks that combine both settings. To fill this gap, Bai et al. (2024) introduced a benchmark specifically designed for federated domain generalization (DG), evaluating methods across diverse datasets with varying levels of client heterogeneity. In our experiments, we use the PACS Li et al. (2017) and iWildCAM data sets from this benchmark to model realistic label shifts between the client and target distributions. We modify the PACS dataset to consist of three clients, each missing a label that is present in the other two. Additionally, one client is reduced to one-tenth the size of the others, and the target distribution is made sparse in the same label as that of the smaller client. Further details are given in Appendix A.
Furthermore, we construct two additional tasks by introducing label shift to standard image classification data sets, Fashion-MNIST (Xiao et al., 2017) and CIFAR-10 (Krizhevsky, 2009). We apply two label shift sampling strategies: sparsity sampling and Dirichlet sampling. Sparsity sampling involves randomly removing a subset of labels from clients and the target domain, following the data set partitioning technique first introduced in McMahan et al. (2017) and extensively used in subsequent studies (Geyer et al., 2017; Li et al., 2020a, 2022). Each client is assigned $C$ random labels, with an equal number of samples for each label and no overlap among clients.
@@
Data set Label split M FedPALS FedAvg FedProx SCAFFOLD Oracle
Fashion-MNIST $C=3$ $10$ $92.4\pm 2.1$ $67.1\pm 22.0$ $66.9\pm 20.8$ $69.5\pm 19.3$ $97.6\pm 2.1$
$C=2$ $80.6\pm 23.7$ $53.9\pm 36.2$ $52.9\pm 35.7$ $54.9\pm 36.8$ $97.5\pm 4.0$
CIFAR-10 $C=3$ $65.6\pm 10.1$ $44.0\pm 8.4$ $43.5\pm 7.2$ $43.3\pm 7.4$ $85.5\pm 5.0$
$C=2$ $10$ $72.8\pm 17.4$ $46.7\pm 15.8$ $47.7\pm 15.6$ $46.7\pm 14.9$ $89.2\pm 3.9$
$\beta=0.1$ $62.6\pm 17.9$ $40.8\pm 9.2$ $41.9\pm 9.7$ $43.5\pm 10.5$ $79.2\pm 3.7$
PACS $C=6$ $3$ $86.0\pm 2.9$ $73.4\pm 1.6$ $75.3\pm 1.3$ $73.9\pm 0.3$ $90.5\pm 0.3$
Dirichlet sampling simulates realistic non-i.i.d. label distributions by, for each client $i$, drawing a sample $p_{i}\sim\mathrm{Dirichlet}_{K}(\beta)$, where $p_{i}(k)$ represents the proportion of samples in client $i$ that have label $k\in[K]$. We use a symmetric concentration parameter $\beta>0$ which controls the sparsity of the client distributions. A smaller $\beta$ results in more heterogeneous client data sets, while a larger value approximates an i.i.d. setting. This widely-used method for sampling clients was first introduced by Yurochkin et al. (2019). While prior works have focused on inter-client distribution shifts assuming that client and target domains are equally distributed, we apply these sampling strategies also to the target set, thereby introducing label shift between the client and target data. Figures 2(b) & 6(b) (latter in appendix) illustrate an example with $C=6$ for sparsity sampling and Dirichlet sampling with $\beta=0.1$, where the last client (Client 9) is chosen as the target. In addition, we investigate the effect of $T(Y)\notin\mathrm{Conv}(S)$ in a synthetic task described in 5.2.
@@
Alongside FedAvg, we use SCAFFOLD and FedProx (Karimireddy et al., 2020; Li et al., 2020b) as baselines, chosen due to their prominence in the literature. SCAFFOLD mitigates client drift in heterogeneous data environments by introducing control variates to correct local updates. FedProx incorporates a proximal term to the objective to limit the divergence of local models from the global model. For the synthetic experiment in Section 5.2, we use a logistic regression model. For CIFAR-10 and Fashion-MNIST, we use small, two-layer convolutional networks, while for PACS and iWildCAM, we use a ResNet-50 pre-trained on ImageNet. Early stopping, model hyperparameters, and $\lambda$ in FedPALS are tuned using a validation set that reflects the target distribution in the synthetic experiment, CIFAR-10, Fashion-MNIST, and PACS. For iWildCAM, we adopt the hyperparameters reported by Bai et al. (2024) and select $\lambda$ using the same validation set used in their work. We report the mean test accuracy and standard deviation for each method over 3 independent random seeds for PACS and iWildCam and 8 seeds for the smaller Fashion-MNIST and CIFAR-10, to ensure a robust evaluation.
@@ 5.1 Experimental results on benchmark tasks
We present summary results for three tasks with selected skews in Table 1 and explore detailed results below. Across these tasks, FedPALS consistently outperforms or matches the best-performing baseline. For PACS, Fashion-MNIST and CIFAR-10, we include results for an Oracle FedAvg model, which is trained on clients whose distributions are identical to the target distribution, eliminating any client-target distribution shift (see Appendix A for details on its construction). A FedPALS Oracle would be equivalent since there is no label shift. The Oracle, which enjoys perfect alignment between client and target distributions, achieves superior performance, underscoring the challenges posed by distribution shifts in real-world scenarios where such alignment is absent.
@@
(c) Accuracy varying labels per client, $C$.
Figure 2: Results on CIFAR-10 with sparsity sampling, varying the number of clients $C$ across 10 clients. Clients with IDs 0–8 are used in training, and Client 9 is the target client. The task is more difficult for small $C$, when fewer clients share labels, and the projection distance is larger.
@@
Figure 3: Target accuracy during training of FedPALS compared to baselines on PACS (a) and iWildCam (b), averaged over 3 random seeds. $M$ is the number of training clients.
@@
CIFAR-10/Fashion-MNIST. Figure 2(c) shows the results for the CIFAR-10 data set, where we vary the label sparsity across clients. In the standard i.i.d. setting, where all labels are present in both the training and target clients ($C=10$), all methods perform comparably. However, as label sparsity increases and fewer labels are available in client data sets (i.e., as $C$ decreases), we observe a performance degradation in standard baselines. In contrast, our proposed method, FedPALS, leverages optimized aggregation to achieve a lower target risk, resulting in improved test accuracy under these challenging conditions. Similar trends are observed for Fashion-MNIST, as shown in Figure 7 in Appendix B. Furthermore, the results in the highly non-i.i.d. cases ($C=2,3$ and $\beta=0.1$) are summarized in Table 1. Additional experiments in Appendix B examine how the algorithms perform with varying numbers of local epochs and clients.
@@
PACS. As shown in Figure 3(a), being faithful to the target distribution is crucial for improved performance. Lower values of $\lambda$ generally correspond to better performance. Notably, FedAvg struggles in this setting because it systematically underweights the client with the distribution most similar to the target, leading to suboptimal model performance. In fact, this even causes performance to degrade over time. Interestingly, the baselines also face challenges on this task: both FedProx and SCAFFOLD perform similarly to FedPALS when $\lambda=93$. However, FedPALS demonstrates significant improvements over these methods, highlighting the effectiveness of our aggregation scheme in enhancing performance. Additionally, in Table 1, we present the models selected based on the source validation set, where FedPALS outperforms all other methods. For comprehensive results, including all FedPALS models and baseline comparisons, please refer to Table 3 in Appendix B.
@@
iWildCam. The test performance across communication rounds is shown in Figure 3(b). Initially, FedPALS widens the performance gap compared to FedAvg, but as training progresses, this gain diminishes. While FedPALS quickly reaches a strong performing model, it eventually plateaus. The rate of convergence and level of performance reached appears to be influenced by the choice of $\lambda$, with lower values of $\lambda$ leading to faster plateaus at lower levels compared to larger ones. This suggests that more uniform client weights and a larger effective sample size are preferable in this task. Given the iWildCam dataset’s significant class imbalance – with many classes having few samples – de-emphasizing certain clients can degrade performance. We also note that our assumption of label shift need not hold in this experiment, as the cameras are in different locations, potentially leading to variations in the conditional distribution $p(X\mid Y)$. The performance of the models selected using the source validation set is shown in Table 2 in Appendix B. There we see that FedPALS performs comparably to FedAvg and FedProx while outperforming SCAFFOLD. Unlike in other tasks, where FedProx performs comparably or worse than FedPALS, we see FedProx achieve the highest F1-score on this task. Therefore, we conduct an additional experiment where we use both FedProx and FedPALS together, as they are not mutually exclusive. This results in the best performing model, see Figure 3(b). Finally, as an illustration of the impact of increasing $\lambda$, we provide the weights of the clients in this experiment alongside the FedAvg weights in 5 in Appendix B. We note that as $\lambda$ increases, the weights increasingly align with those of FedAvg while retaining weight on the clients whose label distributions most resemble that of the target.
5.2 Synthetic experiment: effect of projection distance on test error

### v2

@@ 3.1 FedPALS in the limits
Now, consider when the target label distribution is in the convex hull of the source label distributions, $\mathrm{Conv}(S)$. Then, we can find a convex combination $\alpha^{c}$ of source distributions $S_{i}(Y)$ that recreate $T(Y)$, that is, $T(Y)=\sum_{i=1}^{M}\alpha_{i}^{c}S_{i}(Y)$. However, when there are more clients than labels, $M>K$, such a satisfying combination $\alpha^{c}$ need not be unique and different combinations may have different effective sample size. Let $A^{c}=\{\alpha^{c}\in\Delta^{M-1}:T(Y)=(\alpha^{c})^{\top}S(Y)\}$ denote all satisfying combinations where $S(Y)\in\mathbb{R}^{M\times K}$ is the matrix of all client label marginals. For a sufficiently small regularization penalty $\lambda$, the solution to (9) will be the satisfying combination with largest effective sample size.
@@
If there are fewer clients than labels, $M<K$, the set of target distributions for which a satisfying combination exists has measure zero, see Figure 1 (left). Nevertheless, the two cases above allow us to interpolate between being as faithful as possible to the target label distribution ($\lambda\to 0$) and retaining the largest effective sample size ($\lambda\to\infty$), the latter coinciding with FedAvg. Finally, when $T\in\mathrm{Conv}(S)$ and $\lambda\rightarrow 0$, Proposition 1 applies also to FedPALS; the aggregation strategy results in an unbiased estimate of the target risk gradient in the SGD setting. However, like the unregularized weights, Proposition 1 does not apply for multiple local client updates.
@@
If the target distribution does not lie in $\mathrm{Conv}(S)$, see Figure 1 (left), FedPALS projects the target to the “closest point” in $\mathrm{Conv}(S)$ if $\lambda=0$, and to a tradeoff between this projection and the FedAvg aggregation if $\lambda>0$. We have a discussion on how to choose $\lambda$ in the second and third cases in Appendix A.1.
@@
In problems with a large number of labels, $K\gg 1$, it is common that any individual domain (clients or target) supports only a subset of the labels. For example, in the IWildCam benchmark, not every wildlife camera captures images of all animal species. When the target $T(Y)$ is sparse, meaning $T(y)=0$ for certain labels $y$, it becomes easier to find a good match $(\alpha^{\lambda})^{\top}S(Y)\approx T(Y)$ if the client label distributions are also sparse. Achieving a perfect match, i.e., $T\in\mathrm{Conv}(S)$, requires that (i) the clients collectively cover all labels in the target, and (ii) each client contains only labels that are present in the target. If this is also beneficial for learning, it would suggest that the client-presence of labels that are not present in the target would harm the aggregated model. We study the implications of sparsity of label distributions empirically in Section 5.
@@ 4 Related work
Efforts to mitigate the effects of distributional shifts in federated learning can be broadly categorized into client-side and server-side approaches. Client-side methods use techniques such as clustering clients with similar data distributions and training separate models for each cluster (Ghosh et al., 2020; Sattler et al., 2020; Vardhan et al., 2024), and meta-learning to enable models to quickly adapt to new data distributions with minimal updates (Chen et al., 2018; Jiang et al., 2019; Fallah et al., 2020). Other notable strategies include regularization techniques that penalize large deviations in client updates to ensure stable convergence (Li et al., 2020b, 2021) and recent work on optimizing for flatter minima to enhance model robustness (Qu et al., 2022; Caldarola et al., 2022). Server-side methods focus on improving model aggregation or applying post-aggregation adjustments. These include optimizing aggregation weights (Reddi et al., 2021), learning adaptive weights (Li et al., 2023), iterative moving averages to refine the global model (Zhou et al., 2023), and promoting gradient diversity during updates (Zeng et al., 2023). Both categories of work overlook shifts in the target distribution, leaving this area unexplored.
Another related area is personalized federated learning, which focuses on fine-tuning models to optimize performance on each client’s specific local data (Collins et al., 2022; Boroujeni et al., 2024). This setting differs fundamentally from our work, which focuses on improving generalization to new target clients without any training data available for fine-tuning. Label distribution shifts have also been explored with methods such as logit calibration (Zhang et al., 2022; Xu et al., 2023), novel loss functions (Wang et al., 2021), feature augmentation (Xia et al., 2023), gradient reweighting (Xiao et al., 2023), and contrastive learning (Wu et al., 2023). However, like methods aimed at mitigating the effects of general shifts, these do not address the challenge of aligning models with an unseen target distribution, as required in our setting.
Generalization under domain shift in federated learning remains underdeveloped (Bai et al., 2024). The work most similar to ours is that of agnostic federated learning (AFL) (Mohri et al., 2019), which aims to learn a model that performs robustly across all possible target distributions within the convex hull of client distributions. One notable approach is tailored for medical image segmentation, where clients share data in the frequency domain to achieve better generalization across domains (Liu et al., 2021). However, this technique requires data sharing, making it unsuitable for privacy-sensitive applications like ours. A different line of work focuses on addressing covariate shift in federated learning through importance weighting (Ramezani-Kebrya et al., 2023). Although effective, this method requires sending samples from the test distribution to the server, which violates our privacy constraints.
@@ 5 Experiments
While numerous benchmarks exist for federated learning (Caldas et al., 2018; Chen et al., 2022) and domain generalization (Gulrajani and Lopez-Paz, 2020; Koh et al., 2021), respectively, until recently none have addressed tasks that combine both settings. To fill this gap, Bai et al. (2024) introduced a benchmark specifically designed for federated domain generalization (DG), evaluating methods across diverse datasets with varying levels of client heterogeneity. In our experiments, we use the PACS Li et al. (2017) and iWildCAM data sets from the Bai et al. (2024) benchmark to model realistic label shifts between the client and target distributions. We modify the PACS dataset to consist of three clients, each missing a label that is present in the other two. Additionally, one client is reduced to one-tenth the size of the others, and the target distribution is made sparse in the same label as that of the smaller client. For further details see Appendix A.
Furthermore, we construct two additional tasks by introducing label shift to standard image classification data sets, Fashion-MNIST (Xiao et al., 2017) and CIFAR-10 (Krizhevsky, 2009). We apply two label shift sampling strategies: sparsity sampling and Dirichlet sampling. Sparsity sampling involves randomly removing a subset of labels from clients and the target domain, following the data set partitioning technique first introduced in McMahan et al. (2017). Each client is assigned $C$ random labels, with an equal number of samples for each label and no overlap among clients.
@@
Data set Label split M FedPALS FedAvg FedProx SCAFFOLD AFL FedRS Oracle
Fashion-MNIST $C=3$ 1010 $\mathbf{92.4\pm 2.1}$ $7.1\pm 22.0 66$$6.9\pm 20.8$ $69.5\pm 19.3$ $78.9\pm 14.7$ $85.3\pm 13.5$ $97.6\pm 2.1$
$C=2$ $\mathbf{80.6\pm 23.7}$ $53.9\pm 36.2$ $52.9\pm 35.7$ $54.9\pm 36.8$ $78.6\pm 20.0$ $63.14\pm 20.2$ $97.5\pm 4.0$
CIFAR-10 $C=3$ $\mathbf{65.6\pm 10.1}$ $44.0\pm 8.4$ $43.5\pm 7.2$ $43.3\pm 7.4$ $53.2\pm 0.9$ $44.0\pm 8.0$ $85.5\pm 5.0$
$C=2$ 1010 $\mathbf{72.8\pm 17.4}$ $46.7\pm 15.8$ $47.7\pm 15.6$ $46.7\pm 14.9$ $54.7\pm 0.1$ $49.4\pm 9.5$ $89.2\pm 3.9$
$\beta=0.1$ $\mathbf{62.6\pm 17.9}$ $40.8\pm 9.2$ $41.9\pm 9.7$ $43.5\pm 10.5$ $53.4\pm 11.5$ $57.1\pm 11.2$ $79.2\pm 3.7$
PACS $C=6$ 33 $\mathbf{86.0\pm 2.9}$ $73.4\pm 1.6$ $75.3\pm 1.3$ $73.9\pm 0.3$ $74.5\pm 0.9$ $76.1\pm 1.6$ $90.5\pm 0.3$
Dirichlet sampling simulates realistic non-i.i.d. label distributions by, for each client $i$, drawing a sample $p_{i}\sim\mathrm{Dirichlet}_{K}(\beta)$, where $p_{i}(k)$ represents the proportion of samples in client $i$ that have label $k\in[K]$. We use a symmetric concentration parameter $\beta>0$ which controls the sparsity of the client distributions. See Appendix A.2.
While prior works have focused on inter-client distribution shifts assuming that client and target domains are equally distributed, we apply these sampling strategies also to the target set, thereby introducing label shift between the client and target data. Figures 2(b) & 5(b) (latter in appendix) illustrate an example with $C=6$ for sparsity sampling and Dirichlet sampling with $\beta=0.1$, where the last client (Client 9) is chosen as the target. In addition, we investigate the effect of $T(Y)\notin\mathrm{Conv}(S)$ in a task described in B.4.
@@
Alongside FedAvg, we use SCAFFOLD, FedProx, AFL and FedRS (Karimireddy et al., 2020; Li et al., 2020b; Mohri et al., 2019; Li and Zhan, 2021) as baselines. The first two chosen due to their prominence in the literature for handling non-iid data, and AFL which is similar in concept to FedPALS and aims to optimize for an unseen domain. We also include FedRS, designed specifically to address label distribution skew. For the synthetic experiment in Section B.4, we use a logistic regression model. For CIFAR-10 and Fashion-MNIST, we use small, two-layer convolutional networks, while for PACS and iWildCAM, we use a ResNet-50 pre-trained on ImageNet. Early stopping, model hyperparameters, and $\lambda$ in FedPALS are tuned using a validation set that reflects the target distribution in the synthetic experiment, CIFAR-10, Fashion-MNIST, and PACS. This tuning process consistently resulted in setting the number of local epochs to $E=1$ across all experiments. For iWildCAM, we adopt the hyperparameters reported by Bai et al. (2024) and select $\lambda$ using the same validation set used in their work. We report the mean test accuracy and standard deviation for each method over 3 independent random seeds for PACS and iWildCam and 8 seeds for the smaller Fashion-MNIST and CIFAR-10, to ensure robust evaluation.
@@ 5.1 Experimental results on benchmark tasks
We present results for three tasks with selected skews in Table 1 and explore detailed results below. Across these tasks, FedPALS consistently outperforms or matches the best-performing baseline. For PACS, Fashion-MNIST and CIFAR-10, we include results for an Oracle FedAvg model, which is trained on clients whose distributions are identical to the target distribution, eliminating any client-target distribution shift (see Appendix A for details). A FedPALS-Oracle would be equivalent since there is no label shift. The Oracle, which has perfect alignment between client and target distributions, achieves superior performance, underscoring the challenge posed by distribution shifts in real-world scenarios where such alignment is absent.
@@
(c) Accuracy vs labels per client, $C$.
Figure 2: Results on CIFAR-10 with sparsity sampling, varying the number of labels per clients $C$ across 10 clients. Clients with IDs 0–8 are used in training, and Client 9 is the target client. The task is more difficult for small $C$, when fewer clients share labels, and the projection distance is larger.
@@
Figure 3: Target accuracy/F1-score during training of FedPALS compared to baselines on PACS (a) and iWildCam (b), averaged over 3 random seeds. $M$ is the number of training clients. Non-zero $\lambda$-values chosen to correspond to an ESS of $25\%$, $50\%$ and $75\%$.
@@
CIFAR-10/Fashion-MNIST. Figure 2(c) shows the results for the CIFAR-10 data set, where we vary the label sparsity across clients. In the standard i.i.d. setting, where all labels are present in both the training and target clients ($C=10$), all methods perform comparably. However, as label sparsity increases and fewer labels are available in client data sets (i.e., as $C$ decreases), we observe a performance degradation in standard baselines. In contrast, our proposed method, FedPALS, leverages optimized aggregation to achieve a lower target risk, resulting in improved test accuracy under these challenging conditions. Similar trends are observed for Fashion-MNIST, as shown in Figure 6 in Appendix B. Furthermore, the results in the highly non-i.i.d. cases ($C=2,3$ and $\beta=0.1$) are summarized in Table 1. Additional experiments in Appendix B examine how the algorithms perform with varying numbers of local epochs (up to 40) and clients (up to 100).
@@
PACS. As shown in Figure 3(a), being faithful to the target distribution is crucial for improved performance. Lower values of $\lambda$ generally correspond to better performance. Notably, FedAvg struggles in this setting because it systematically underweights the client with the distribution most similar to the target, leading to suboptimal model performance. In fact, this even causes performance to degrade over time. Interestingly, the baselines also face challenges on this task: both FedProx, FedRS and SCAFFOLD perform similarly to FedPALS when $\lambda=93$. However, FedPALS demonstrates significant improvements over these methods, highlighting the effectiveness of our aggregation scheme in enhancing performance. We also see that FedPALS + FedProx performs comparably to just using FedPALS in this case, although it does have higher variance. Additionally, in Table 1, we present the models selected based on the source validation set, where FedPALS outperforms all other methods. For comprehensive results, including all FedPALS models and baseline comparisons, refer to Table 4 in Appendix B.
@@
iWildCam. The test performance across communication rounds is shown in Figure 3(b). Initially, FedPALS widens the performance gap compared to FedAvg, but as training progresses, this gain diminishes. While FedPALS quickly reaches a strong performing model, it eventually plateaus. The rate of convergence and level of performance reached appears to be influenced by the choice of $\lambda$, with lower values of $\lambda$ leading to faster plateaus at lower levels compared to larger ones. This suggests that more uniform client weights and a larger effective sample size are preferable in this task. Given the iWildCam dataset’s significant class imbalance – with many classes having few samples – de-emphasizing certain clients can degrade performance. We also note that our assumption of label shift need not hold in this experiment, as the cameras are in different locations, potentially leading to variations in the conditional distribution $p(X\mid Y)$. The performance of the models selected using the source validation set is shown in Table 3 in Appendix B. There we see that FedPALS performs comparably to FedAvg and FedProx while outperforming SCAFFOLD. Unlike in other tasks, where FedProx performs comparably or worse than FedPALS, we see FedProx achieve the highest F1-score on this task. Therefore, we conduct an additional experiment where we use both FedProx and FedPALS together, as they are not mutually exclusive. This results in the best performing model, see Figure 3(b). Due to memory issues with the implementation FedRS was not able to run for this experiment and is omitted. AFL fails to learn in this task and is thus also omitted, although results are shown in Table 3 in Appendix B. Finally, as an illustration of the impact of increasing $\lambda$, we provide the weights of the clients in this experiment alongside the FedAvg weights in 4 in Appendix B. We note that as $\lambda$ increases, the weights increasingly align with those of FedAvg while retaining weight on the clients whose label distributions most resemble that of the target.